OpenStudy (anonymous):

A pet shop had 8 identical fish tanks in the shape of rectangular prisms. The width and height of each tank was (x - 7) inches and the length was 8 inches more than its width. The expression below shows the maximum amount of water which all the tanks together can hold. 8(x + 1)(x - 7)2 What does the factor 8(x + 1) represent? The total height of all the tanks The total length of all the tanks The total volume of all the tanks The total base area of all the tanks

OpenStudy (anonymous):

@hyntc

OpenStudy (anonymous):

The volume of all the tanks 8 identical tanks Volume of each tank; L*W*H = (x-7+8)(x-7)(x-7) = (x+1)(x-7)^2 so, for 8 tanks: 8(x+1)(x-7)^2